OpenStudy (anonymous):

Suppose you are playing a game with two number cubes. Let A represent rolling 2, 3, or 4, and B represent rolling 1, 5, or 6. The probability of A is 1/2 and the probability of B is 1/2. a. Simplify b. What is the probability that one number cube shows 2, 3, or 4, and the other shows 1, 5, or 6? Treat part a as a binomial that is squared. For part b, Don’t forget to think about all the ways the desired events could occur

OpenStudy (anonymous):

ohh....i dont do probability sorry :\

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

since two cubes are being used, let the sample space be 6^2 =36

OpenStudy (anonymous):

and since P(A)=1/2 and P(B)=1/2 then for second bit we should find p(AandB) = p(A) + p(B) - p(AUB) = 1/2 +1/2 - p(AUB)

OpenStudy (anonymous):

ok ok i see what i did wrong i put the answer as 1/4, and 1/4